OpenStudy (anonymous):

Solve the system using elimination, explain your steps 2x + 3y = 5 8y = x + 26

OpenStudy (anonymous):

@tHe_FiZiCx99 can u help me please

OpenStudy (anonymous):

@Mertsj

OpenStudy (anonymous):

@kelliegirl33 @kelly226

OpenStudy (the_fizicx99):

It would be alot easier to use the elimination method here to isolate y and then plug that tin but ok.. 2x + 3y = 5 8y = x + 26 -x -x 2x + 3y = 5 -x + 8y = 26 Multiply the top by -1 and the bottom by -2.  -2x -3y = -5 2x - 16y = -52 Now add them, -2x + 2x cancel out. -3y + (-16y) = -19y -19y = -57 /-19 /-19 y = 3 Now plug in 3 for y and solve for x.